OpenStudy (help!!!!):

Some help me out this is ALG 1 problem. Steps are appreciated

OpenStudy (help!!!!):

when two numbers are added together, the result is 45. Twice the greater number is 6 more than 5 times the lesser. What are the numbers?

OpenStudy (anonymous):

Number 1 = x Number 2 = y ---------- x + y = 45 (let y be greater, it doesn't matter as long as only one is greater) 2y = 5x +6

OpenStudy (anonymous):

Solve the system

OpenStudy (anonymous):

Do you know how?

OpenStudy (help!!!!):

wait does this work?

OpenStudy (help!!!!):

x+y=45 2x+6+5y=45 x=greater y=lesser

OpenStudy (help!!!!):

wait nevermind

OpenStudy (help!!!!):

can you continue in your way?

OpenStudy (anonymous):

If x is lesser, than just switch the x and y in the second equation I gave you

OpenStudy (anonymous):

*greater

OpenStudy (help!!!!):

yeah, could you explain your way?

OpenStudy (anonymous):

2x = 5y + 6 divide each side by 2 x = (5/2)y+3 ------------ x + y = 45 (5/2)y+3 + y = 45 (5/2)y + y + 3 = 45 (7/2)y +3 = 45 -3 on both sides (7/2)y = 42 divide both sides by (7/2) y = 12 --------- x + 12 = 45 subtract 12 on both sides x = 33 ------- (33, 12)

OpenStudy (help!!!!):

Thank you so much :DDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

Just substitute x in the terms of y for y to solve for y, then use that to solve for x